A motorboat is observed to travel10kmph relative to earth in the direction 37deg north of east.If the velocity of boat due to the wind only is 2kmph westward and that due to current only is 4kmph southward,what is the magnitude and direction  of the velocity of the boat due to its own power.

Grade:Upto college level

2 Answers

AskIITians Expert PRASAD IIT Kharagpur
14 years ago

East -  i vector  North - j vector.

Vel of boat = vel of itself  + vel of wind + vel of current.

Vb = Vit + Vw + Vr

10(cos 37 + sin 37 j) = Vit  + (-2 i) + (-4 j)

8 i + 6 j = Vit  + (-2 i) + (-4 j)

Vit = 10 i + 10 j

Boat travels at 14.14 kmph 45deg north of east due to its owm power.

 

Ramesh V
14 years ago

vel. of boat w.r.t wind Vbw = -2j

vel. of boat w.r.t water current Vbc = -4i

given vel. of boat w.r.t earth(water current + wind): Vb(w+c) = 6i + 8j

so, Vb - Vw = -2j                   .............(1)

       Vb - Vc = -4i                   .............(2)

       Vb - V(w+c) = 6i + 8j     . ............(3)

Let vel. of boat alone due to its power is = (ai+bj)

From above (1) + (2) -(3) gives

Vb = -10.(i + j)

The magintude of vel. of boat = 14.14 kmph

direction is 45 degrees south west

---

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch